Errata for Volume minimizing submanifolds in compact symmetric spaces

(Kodai Math. J., vol.17 no.3 (1994), 456–462)

Hiroyuki Tasaki

April 12, 2007

p.456 Line 9: “othor” should read “other”
p.456 Line 6 from bottom: “clibration” should read “calibration”
p.456 Line 5 from bottom: “Wirtnger’s” should read “Wirtinger’s”
p.456 Line 1 from bottom: “any” should read “and”
p.459 Line 9 from bottom: “tnagent” should read “tangent”
p.460 Line 7 from bottom: “summetric” should read “symmetric”